OpenStudy (anonymous):

which equation will have the same solution as 3-5(2x+1=3(3+x)?

OpenStudy (anonymous):

Solve for x: 2 x+1 = 3 (x+3) Move everything to the left hand side. Subtract 3 (x+3) from both sides of 2 x+1 = 3 (x+3): 1+2 x-3 (x+3) = 3 (x+3)-3 (x+3) Look for two terms that sum to zero. 3 (x+3)-3 (x+3) = 0: 1+2 x-3 (x+3) = 0 Distribute -3 over x+3. -3 (x+3) = -3 x-9: 1+2 x+-9-3 x = 0 Group like terms in 1+2 x-9-3 x. Grouping like terms, 1+2 x-9-3 x = (2 x-3 x)+(1-9): (2 x-3 x)+(1-9) = 0 Combine like terms in 2 x-3 x. 2 x-3 x = -x: -x+(1-9) = 0 Evaluate 1-9. 1-9 = -8: -x+-8 = 0 Isolate terms with x to the left hand side. Add 8 to both sides: -x+(-8+8) = 8 Look for two terms that sum to zero. -8+8 = 0: -x = 8 Solve for x. Multiply both sides of -x = 8 by -1: -(x)/(-1) = -8 Any nonzero number divided by itself is one. (-1)/(-1) = 1: Answer: | | x = -8 So another equation with the same answer would be -x-8 = 0 right?